OpenStudy (anonymous):

Two beauticians, Lydia and Jackie, are working on your hair. Lydia can complete the work in 2 hours, while Jackie can complete the work in 1 hour. How many hours do the beauticians take if they work together?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

I also need this one.

OpenStudy (phi):

You could think of these problems as rate * time = job rate is "jobs per hour", and time is hours Lydia can complete the work in 2 hours... that means Lydia's rate is 1 job/2 hours

OpenStudy (anonymous):

The choices for the first question are: -1/2 -3/2 -2/3 -4/3

OpenStudy (anonymous):

I tried to solve it but I got 1...

OpenStudy (potatoes.ramu):

I don't know how to do this, I'm so sorry :(

OpenStudy (phi):

Jackie can complete the work in 1 hour means Jackie's rate is 1 job per 1 hour or 1 job/hour

OpenStudy (anonymous):

so 1/2 + 1/1

OpenStudy (anonymous):

1/2 + 1/1= 1/x

OpenStudy (anonymous):

I know *how* to do it. it's just that I got a whole number instead of a fraction

OpenStudy (phi):

If you do rate * time for both of them: \[\left( \frac{1}{2}+ 1\right) t = 1 \text{ job}\]

OpenStudy (anonymous):